Three small positive electrical charges A, B and C are placed along a straight line. The distance between A and B is 12 cm. A and B have equal charges while C has twice the charge of each of A and B.



At what distance from B should C be placed in order for the resulting force on B to be zero?


1
Expert's answer
2021-01-27T07:37:19-0500

Solution:


A and B have charge q, C has charge 2q, correct? Assume that is correct.

A's field at B points right. C's field at B points left so the fields subtract.

"\dfrac{q}{0.12^2}-\dfrac{2q}{r^2}=0" where r is the distance from C to B


"q*r^2=2q*0.12^2"

"r^2=2*0.12^2 =0.288"

r = "\sqrt{\smash[b]{0.288}}" = 0.17m  = 17cm from B to C